69. Suppose a monopolistically competitive market has 10 firms. The largest firm has a 90 percent share of the market and the other nine firms each have 1 percent of the market. The Herfindahl-Hirschman Index for this industry is ________.

a. 8,109

b. 8,100

c. 99

d. 909

70. An industry is deemed concentrated when ________.

a. each firm in that industry has a small market share

b. all the firms in that industry charge a price lower than the average cost of production

c. a few firms account for a large fraction of the total sales in that industry

d. most of the firms in that industry earn zero economic profits in the long run

71. Which of the following markets is likely to be a monopoly?

a. The market for shower gel

b. The market for smart phones

c. The market for wheat

d. The market for patented nuclear medicines

72. Which of the following markets is likely to be perfectly competitive?

a. The market for shower gel

b. The market for wheat

c. The market for patented nuclear medicines

d. The market for smart phones
Answer #1

69) HHI = 90^2+1^2*9 = 8100+9 = 8109

Option(A)

70) a few firms account for a large fraction of total sales in that industry

option(C)

71) d. The market for patented nuclear medicines

72) b. The market for wheat